# Depth Anything: Unleashing the Power of Large-Scale Unlabeled Data

### 19 Jan 2024 v1

@@ Abstract
The work was done during an internship at TikTok.
@@ 1 Introduction
The field of computer vision and natural language processing is currently experiencing a revolution with the emergence of “foundation models” [6] that demonstrate strong zero-/few-shot performance in various downstream scenarios [44, 58]. These successes primarily rely on large-scale training data that can effectively cover the data distribution. Monocular Depth Estimation (MDE), which is a fundamental problem with broad applications in robotics [65], autonomous driving [63, 79], virtual reality [47], etc., also requires a foundation model to estimate depth information from a single image. However, this has been underexplored due to the difficulty of building datasets with tens of millions of depth labels. MiDaS [45] made a pioneering study along this direction by training an MDE model on a collection of mixed labeled datasets. Despite demonstrating a certain level of zero-shot ability, MiDaS is limited by its data coverage, thus suffering disastrous performance in some scenarios.
@@ 4.1 Implementation Details
We adopt the DINOv2 encoder [42] for feature extraction. Following MiDaS [45, 5], we use the DPT [46] decoder for depth regression. All labeled datasets are simply combined together without re-sampling. In the first stage, we train a teacher model on labeled images for 20 epochs. In the second stage of joint training, we train a student model to sweep across all unlabeled images for one time. The unlabeled images are annotated by a best-performed teacher model with a ViT-L encoder. The ratio of labeled and unlabeled images is set as 1:2 in each batch. In both stages, the base learning rate of the pre-trained encoder is set as 5e-6, while the randomly initialized decoder uses a 10$\times$ larger learning rate. We use the AdamW optimizer and decay the learning rate with a linear schedule. We only apply horizontal flipping as our data augmentation for labeled images. The tolerance margin $\alpha$ for feature alignment loss is set as 0.15. For more details, please refer to our appendix.
@@ 4.5 Ablation Studies
Comparison with DINOv2 in downstream tasks. We have demonstrated the superiority of our trained encoder when fine-tuned to downstream tasks. Since our finally produced encoder (from large-scale MDE training) is fine-tuned from DINOv2 [42], we compare our encoder with the original DINOv2 encoder in Table 11. It can be observed that our encoder performs better than the original DINOv2 encoder in both the downstream metric depth estimation task and semantic segmentation task. Although the DINOv2 weight has provided a very strong initialization (also much better than the MiDaS encoder as reported in Table 10), our large-scale and high-quality MDE training can further enhance it impressively in downstream transferring performance.
Figure 4: We compare our depth prediction with MiDaS. Meantime, we use ControlNet to synthesize new images from the depth map (the last row). First row: input image, second row: depth prediction.
@@ 4.6 Qualitative Results
We visualize our model predictions on the six unseen datasets in Figure 3. Our model is robust to test images from various domains. In addition, we compare our model with MiDaS in Figure 4. We also attempt to synthesis new images conditioned on the predicted depth maps with ControlNet [84]. Our model produces more accurate depth estimation than MiDaS, as well as better synthesis results, although the ControlNet is trained with MiDaS depth. For more accurate synthesis, we have also re-trained a better depth-conditioned ControlNet based on our Depth Anything, aiming to provide better control signals for image synthesis and video editing. Please refer to our project page or the following supplementary material for more qualitative results,
@@ 7 More Ablation Studies
The necessity of tolerance margin for feature alignment. As shown in Table 12, the gap between the tolerance margin of 0 and 0.15 or 0.30 clearly demonstrates the necessity of this design (mean AbsRel: 0.188 vs. 0.175).

### 07 Apr 2024 v2

@@ Abstract
Work was done during an internship at TikTok.
@@ 1 Introduction
The field of computer vision and natural language processing is currently experiencing a revolution with the emergence of “foundation models" [6] that demonstrate strong zero-/few-shot performance in various downstream scenarios [45, 59]. These successes primarily rely on large-scale training data that can effectively cover the data distribution. Monocular Depth Estimation (MDE), which is a fundamental problem with broad applications in robotics [66], autonomous driving [64, 80], virtual reality [48], etc., also requires a foundation model to estimate depth information from a single image. However, this has been underexplored due to the difficulty of building datasets with tens of millions of depth labels. MiDaS [46] made a pioneering study along this direction by training an MDE model on a collection of mixed labeled datasets. Despite demonstrating a certain level of zero-shot ability, MiDaS is limited by its data coverage, thus suffering disastrous performance in some scenarios.
@@ 4.1 Implementation Details
We adopt the DINOv2 encoder [43] for feature extraction. Following MiDaS [46, 5], we use the DPT [47] decoder for depth regression. All labeled datasets are simply combined together without re-sampling. In the first stage, we train a teacher model on labeled images for 20 epochs. In the second stage of joint training, we train a student model to sweep across all unlabeled images for one time. The unlabeled images are annotated by a best-performed teacher model with a ViT-L encoder. The ratio of labeled and unlabeled images is set as 1:2 in each batch. In both stages, the base learning rate of the pre-trained encoder is set as 5e-6, while the randomly initialized decoder uses a 10$\times$ larger learning rate. We use the AdamW optimizer and decay the learning rate with a linear schedule. We only apply horizontal flipping as our data augmentation for labeled images. The tolerance margin $\alpha$ for feature alignment loss is set as 0.85. For more details, please refer to our appendix.
@@ 4.5 Ablation Studies
Comparison with DINOv2 in downstream tasks. We have demonstrated the superiority of our trained encoder when fine-tuned to downstream tasks. Since our finally produced encoder (from large-scale MDE training) is fine-tuned from DINOv2 [43], we compare our encoder with the original DINOv2 encoder in Table 11. It can be observed that our encoder performs better than the original DINOv2 encoder in both the downstream metric depth estimation task and semantic segmentation task. Although the DINOv2 weight has provided a very strong initialization, our large-scale and high-quality MDE training can further enhance it impressively in downstream transferring performance.
Figure 4: We compare our depth prediction with MiDaS. Meantime, we use ControlNet to synthesize new images from the depth map.
@@ 4.6 Qualitative Results
We visualize our model predictions on the six unseen datasets in Figure 3. Our model is robust to test images from various domains. In addition, we compare our model with MiDaS in Figure 4. We also attempt to synthesis new images conditioned on the predicted depth maps with ControlNet [85]. Our model produces more accurate depth estimation than MiDaS, as well as better synthesis results. For more accurate synthesis, we re-trained a better depth-conditioned ControlNet based on our Depth Anything, aiming to provide better control signals for image synthesis and video editing. Please refer to our project page for more qualitative results on video editing [35] with our Depth Anything.
@@ 5 Conclusion
Acknowledgement. This work is supported by the National Natural Science Foundation of China (No. 62201484), HKU Startup Fund, and HKU Seed Fund for Basic Research.
@@ 7 More Ablation Studies
The necessity of tolerance margin for feature alignment. As shown in Table 12, the gap between the tolerance margin of 1.00 and 0.85 or 0.70 clearly demonstrates the necessity of this design (mean AbsRel: 0.188 vs. 0.175).